Constants Part A The energy flow to the earth from sunlight is about 1.40 kW/m2 Find the maximum value of the electric field for a sinusoidal wave of this intensity N/C Submit Part B Find the maxiiurm value of the magnei field or a sinusoidal wave of this inltensily. Submit RequestAnswer Part C The distance from the earth to the sun is about 1.50x101. Find the total power radiated by the sun Submit Request Answer <Return to Assignment Provide Feedback
Answer #1

a) Emax = sqrt [2 I / e0 c]

= sqrt [2 * 1.4 * 103 / (3 * 108 * 8.85 * 10-12)]

maximum electric field = 1.026 * 103 N/C

b) Bmax = Emax / c

= (1.026 * 103) / (3 * 108)

maximum magnetic field = 3.42 * 10-6 T

c) P = I A = I * 4 pi r2

= 1.4 * 103 * 4 pi * (1.5 * 1011)2

= 3.96 * 1026 W